# Exploring The Missing Semantics In Event Modality

### v1

@@ Abstract
Event cameras offer distinct advantages such as low latency, high dynamic range, and efficient motion capture. However, event-to-video reconstruction (E2V), a fundamental event-based vision task, remains challenging, particularly for reconstructing and recovering semantic information. This is primarily due to the nature of the event camera, as it only captures intensity changes, ignoring static objects and backgrounds, resulting in a lack of semantic information in captured event modality. Further, semantic information plays a crucial role in video and frame reconstruction, yet is often overlooked by existing E2V approaches. To bridge this gap, we propose Semantic-E2VID, an E2V framework that explores the missing visual semantic knowledge in event modality and leverages it to enhance event-to-video reconstruction. Specifically, Semantic-E2VID introduces a cross-modal feature alignment (CFA) module to transfer the robust visual semantics from a frame-based vision foundation model, the Segment Anything Model (SAM), to the event encoder, while aligning the high-level features from distinct modalities. To better utilize the learned semantic feature, we further propose a semantic-aware feature fusion (SFF) block to integrate learned semantics in frame modality to form event representations with rich semantics that can be decoded by the event decoder. Further, to facilitate the reconstruction of semantic information, we propose a novel Semantic Perceptual E2V Supervision that helps the model to reconstruct semantic details by leveraging SAM-generated categorical labels. Extensive experiments demonstrate that Semantic-E2VID significantly enhances frame quality, outperforming state-of-the-art E2V methods across multiple benchmarks. The sample code is included in the supplementary material.
I Introduction
Figure 1: (a): A challenge of event-to-video reconstruction is: event camera only captures intensity changes, containing few semantic information, (b): video reconstruction task requires recovering the missing semantics. Thus, (c): using only events that contain few semantics as input leads to problems such as lacking semantic details. Our approach explores the missing semantics in event modality and utilizes them to produce more accurate results.
@@ I Introduction
Inspired by the human retina, neuromorphic event sensors, to overcome conventional frame-based cameras’ inherent limitations, output a new modality called events, which, unlike standard camera frames that capture image intensities at fixed frame rates, event streams, triggered asynchronously, respond to changes in scene brightness at the pixel level [1, 2]. Specifically, each pixel independently detects and records brightness variations in a log scale, generating an event whenever the change surpasses a predefined threshold. This unique mechanism enables event cameras to achieve ultra-low latency (on the order of microseconds), low power consumption, and an exceptionally high dynamic range (HDR), making them particularly advantageous for capturing fast-moving scenes and extreme lighting conditions [3, 4, 5, 6, 7, 8].
To better interpret event modality, one of the fundamental tasks in event-based vision is event-to-video (E2V) reconstruction, which aims to recover intensity videos from raw event streams. Recent approaches [3, 9, 10, 11, 12] have explored learning-based methods to reconstruct frames in a recurrent encoder-decoder fashion where an event encoder is used to extract deep features and a decoder is used to reconstruct frame results. However, most existing E2V approaches ignore an important ill-posed problem: captured event modality naturally loses valuable semantic knowledge such as object categories, texture, and contextual cues, compared to frame-based intensity modality [1], as shown in Fig. 1 (a). This is because event cameras only capture intensity changes, containing only edge-like structures, and ignore static background information, resulting in a lack of semantics, including category, context relationships, etc, in those regions. Yet, semantic information is vital in video reconstruction and scene understanding as it can provide valuable context for learning and generating high-quality reconstructions [13, 14, 15, 16], shown in Fig. 1 (b). Therefore, a solution is needed to bridge this research gap.
To explore the missing semantics in event modality, an intuitive approach is to use vision foundation models such as the Segment Anything Model [17] (SAM) to distill rich semantic knowledge from frame modality [18, 19]. However, several technical challenges exist. i): a modality mismatch problem hinders naively distilling knowledge from SAM’s frame encoder: in the current E2V pipeline, encoders are used to encode features from input event modality, rather than frame modality. The encoded student features from the event encoder and teacher features extracted from the ground truth frame by SAM’s frame encoder will be misaligned, and the direct distillation between two distinct modalities will heavily affect the event encoder learning and training, leading to poor decoding and reconstruction results. ii): though given that the frame feature with rich semantics has been successfully distilled and obtained, how to incorporate the semantic knowledge of the distilled student feature into the encoded event feature remains unsolved: a solution is needed to first extract the semantics, then fuse them into event features with rich semantics for later decoders to reconstruct accurate videos. iii): existing E2V pipelines either focus on pixel-level [20] supervision or perpetual supervision [3, 10], while ignoring semantic-level supervision. This may result in reconstructions that are visually plausible at the pixel or texture level but lack high-level semantic consistency, such as object boundaries or structure preservation. In all, these challenges are non-trivial, and without explicitly modeling semantic information, the E2V reconstructions may fail to capture important scene understanding cues.
To overcome these challenges, we propose Semantic-E2VID, a framework that explores the missing semantic knowledge and leverages it to improve E2V tasks. Specifically, to transfer robust semantics from a frame modality into the event modality and overcome the misalignment problem triggered across different modalities, we introduce a Cross-Modal Feature Alignment (CFA) module. The CFA module helps to align the high-level features encoded by the event encoder with features encoded by SAM and produce rich semantics that are overseen by event modality without disrupting the event encoding and learning process. To further utilize the learned semantics, we propose a Semantic-aware Feature Fusion (SFF) block, which integrates the semantic knowledge produced by CFA into the encoded event feature representations. This produces event features rich in semantics that are subsequently decoded into high-quality frames. Moreover, to better guide the semantic reconstruction process, we propose a novel Semantic Perceptual E2V Supervision that is usable for any E2V method. By leveraging SAM-generated high-level categorical labels, our supervision strategy encourages the model to preserve attention on rich semantic areas, such as within object boundaries during frame reconstruction, leading to improved reconstruction fidelity. Extensive experiments are done on a total of six public benchmarks, including three datasets with referenced ground truth frames, two datasets with non-referenced frames, and a colorful reconstruction dataset. Semantic-E2VID achieves superior performance compared to previous state-of-the-art E2V approaches across multiple benchmarks and metrics, both quantitatively and qualitatively. The outlined key technical contributions are:
  • We present Semantic-E2VID, the first visual semantic-aware E2V framework that addresses the inherent lack of visual semantic information problem in event modality and improves event-to-video reconstruction quality.
  • We introduce a Cross-Modal Feature Alignment (CFA) module that transfers robust semantics from SAM to the event encoder, while not interpreting the original feature learning process in event modality.
  • We propose a Semantic-aware Feature Fusion (SFF) block that effectively integrates learned semantic features from the frame modality into event feature representations, thereby enhancing frame decoding.
  • We develop a novel Semantic Perceptual E2V Supervision for general E2V tasks and methods, which leverages SAM-generated high-level categorical labels to guide the model toward reconstructing semantically consistent frames with accurate semantic information.
II Related Work
@@
Event-To-Video Reconstruction: Event cameras capture scene dynamics by recording asynchronous brightness changes, offering advantages such as low latency, high dynamic range, and energy efficiency. One fundamental task in event-based vision is event-to-video (E2V) reconstruction, where the goal is to synthesize intensity frames from raw event streams. works [3, 9, 21, 10, 11, 22, 23] have been done using either a recurrent-based, transformer-based, or Mamba-based network to learn event features over time for frame synthesis. Some [24, 25] also explored HDR frame reconstruction. Despite these advancements, most existing E2V methods overlook semantic information such as object categories, texture, relationships, positions, etc, which are missing in the event modality yet important for video reconstruction. [26] leverages text description for semantic-aware E2V, but introduces text, as an extra input modality. Our proposed Semantic-E2VID is the first work that addresses this gap by distilling visual semantic knowledge from SAM to improve semantic learning in event modality and reconstruction quality.
@@
Learning Semantics From SAM: The Segment Anything Model (SAM) [17] has emerged as a powerful vision foundation model, demonstrating strong generalization capabilities across diverse vision tasks [27, 28, 29]. SAM’s success is attributed to its extensive pretraining on large-scale datasets, enabling it to produce robust and transferable semantic representations. Due to its strong semantic understanding, researchers have explored SAM for different downstream tasks. Works [30, 31] have first been done to transfer knowledge from SAM to medical image segmentation. Beyond that, some have utilized SAM as a teacher model for semantic knowledge distillation on other computer vision tasks, such as 2D Image Restoration [32] and feature learning and matching [33]. [34] distill knowledge from SAM for event-based segmentation. However, since both the teacher and student operate on the same task (segmentation), the distillation can be well aligned. In contrast, for the E2V task, directly distilling from frame-based models is challenging in transferring semantic knowledge from the frame modality to the event modality, as the objectives and data representations differ significantly [35].
III Methodology
III-A Modality and Task Formulation
@@
where $E$ represents the recorded events, while $I_{t}$ and $I_{t-w}$ denote the intensity images at two different timestamps with a time window $w$. The function $\Gamma\left\{\theta,\epsilon\right\}$ maps log-intensity changes to discrete event outputs, with $c$ serving as an offset to avoid $\log(0)$. The parameter $\epsilon$ defines the event triggering threshold. Specifically, $\Gamma\{\theta,\epsilon\}=1$ when $\theta\geq\epsilon$, meaning a positive event triggered, and $\Gamma\{\theta,\epsilon\}=-1$ when $\theta\leq-\epsilon$, meaning a negative event triggered [36, 37].
@@
III-B Building From Scratch
We start by implementing a baseline following previous works [3, 10] by adopting a U-Net-based recursive encoder-decoder framework from [3], and including a context fusion hypernetwork module (CFHM) from [11]. Specifically, each encoder module contains a convolutional layer and a ConvLSTM, where the kernel size of the convolutional layer is $5$ and the stride is $2$, while the ConvLSTM has a kernel size of $3$ to maintain consistency in the spatial and channel dimensions of its input and internal state. Each residual block is composed of two convolutional layers, both with a kernel size of $3$. Given an input event voxel grid $V_{k}\in\mathbb{R}^{W\times H\times B}$, the encoder will extract it into event feature $\mathcal{F}_{e}\in\mathbb{R}^{\frac{1}{2}W\times\frac{1}{2}H\times 256}$. For decoder layers, it is composed of a bilinear upsampling and a convolutional layer, and it decodes the high-level feature into the final video frames. Additionally, the CFHM generates dynamic convolutional weights by combining the current event voxel grid with the content information of the previously reconstructed frames, which helps the decoder to generate coherent video.
Although previous studies have made significant progress in the E2V task, the lack of semantics challenge in event modality has been overlooked and prevented previous works from perceiving the semantic information of the scene. Thus, we propose Semantic-E2VID to fully explore semantics in event modality and further aid the E2V task.
@@ III-B Building From Scratch
Figure 2: Overview of our proposed framework. Given input events, we extract event features $\mathcal{F}_{e}$ using an encoder. To learn rich semantics from SAM without misleading the event encoder, the Cross-Modal Feature Alignment (CFA) module aligns $\mathcal{F}_{e}$ with semantics $\mathcal{F}_{sam}$, producing learned semantics $\mathcal{F}_{semantic}$. The Semantic-aware Feature Fusion (SFF) block further enhances $\mathcal{F}_{e}$ by leveraging $\mathcal{F}_{semantic}$, yielding event features with rich semantics $\mathcal{F}_{e}^{\prime}$. A decoder reconstructs the final frame supervised by reconstruction and distillation losses.
III-C Semantic-E2VID
Semantic-E2VID aims to address the absence of semantic information in event data and enhance event-to-video reconstruction. As discussed, several challenges arise: i), event data inherently contains limited semantic cues, necessitating a method to extract and learn rich semantics while correctly aligning high-level features across different modalities; ii), once the semantics in the training frames are learned, an effective fusion strategy is required to integrate them with the encoded event features, producing event features with enriched semantics that facilitate decoders for accurate video reconstruction; and iii), additional supervision that is beneficial to guide the model toward reconstructing frames with semantically meaningful content is highly desired.
To tackle these challenges, Semantic-E2VID first introduces the Cross-Modal Feature Alignment (CFA) module to extract and learn rich semantics from SAM, preventing feature misalignment and misleading. Then, the Semantic-aware Feature Fusion (SFF) block fuses the learned semantics with the encoded event features to generate semantic event features for decoding. Moreover, Semantic Perpetual E2V Supervision is employed to better supervise the reconstruction process, ensuring that the generated frames preserve high-level semantic knowledge. Each component is detailed in the following subsections.

### v2

# Semantic-E2VID: a Semantic-Enriched Paradigm for Event-to-Video Reconstruction

@@ Abstract
Event cameras provide a promising sensing modality for high-speed and high-dynamic-range vision by asynchronously capturing brightness changes. A fundamental task in event-based vision is event-to-video (E2V) reconstruction, which aims to recover intensity videos from event streams. Most existing E2V approaches formulate reconstruction as a temporal–spatial signal recovery problem, relying on temporal aggregation and spatial feature learning to infer intensity frames. While effective to some extent, this formulation overlooks a critical limitation of event data: due to the change-driven sensing mechanism, event streams are inherently semantically under-determined, lacking object-level structure and contextual information that are essential for faithful reconstruction. In this work, we revisit E2V from a semantic perspective and argue that effective reconstruction requires going beyond temporal and spatial modeling to explicitly account for missing semantic information. Based on this insight, we propose Semantic-E2VID, a semantic-enriched end-to-end E2V framework that reformulates reconstruction as a process of semantic learning, fusing and decoding. Our approach first performs semantic abstraction by bridging event representations with semantics extracted from a pretrained Segment Anything Model (SAM), while avoiding modality-induced feature drift. The learned semantics are then fused into the event latent space in a representation-compatible manner, enabling event features to capture object-level structure and contextual cues. Furthermore, semantic-aware supervision is introduced to explicitly guide the reconstruction process toward semantically meaningful regions, complementing conventional pixel-level and temporal objectives. Extensive experiments on six public benchmarks demonstrate that Semantic-E2VID consistently outperforms state-of-the-art E2V methods. By explicitly modeling and leveraging semantic information, our framework produces sharper reconstructions with clearer object boundaries, stronger semantic consistency, and improved robustness under sparse or ambiguous event conditions.
keywords:
Event Camera , Event-based Vision , Event-based Video Reconstruction.
journal: Pattern Recognition
[…]
[1]organization=Department of Electrical and Electronic Engineering,
[…]
School of Computer Science, Peking University,city=Beijing, postcode=100871, country=China thanks: Corresponding author: elam@eee.hku.hk
{graphicalabstract}[Uncaptioned image]
{highlights}
Reformulates event-to-video reconstruction from a semantic viewpoint
Proposes a new E2V paradigm with Semantic Abstraction, Semantic Feature Fusion, and Semantic-Aware Supervision
Extracts semantic knowledge to alleviate the semantic under-determinacy of event data
Achieves structurally and semantically consistent reconstruction with SOTA accuracy on six benchmarks
1 Introduction
Figure 1: (a) Events capture only sparse brightness changes and inherently lack rich semantic information. (b) Traditional E2V models primarily rely on temporal and spatial modeling through feature encoding and decoding, which leads to features with weak semantic awareness. (d) As a result, reconstructed frames often exhibit a blurry appearance and ambiguous object structures. (c) In contrast, our paradigm explicitly incorporates semantic abstraction, semantic feature fusion, and semantic-aware supervision. Semantic knowledge is extracted from the pretrained SAM model and integrated into the latent space to form event deep features with rich semantics. (e) This paradigm enables the reconstruction of sharp frames with stronger semantic consistency, clearer object boundaries, and more faithful structural details.
@@ 1 Introduction
Inspired by the human retina, neuromorphic event sensors are designed to overcome the inherent limitations of conventional frame-based cameras by outputting a novel sensing modality known as events. Unlike standard cameras that capture image intensities at fixed frame rates, event cameras operate asynchronously and respond only to changes in scene brightness at the pixel level [9, 4, 16]. Specifically, each pixel independently detects brightness variations in the logarithmic domain and triggers an event whenever the change exceeds a predefined threshold. This sensing mechanism enables event cameras to achieve ultra-low latency (on the order of microseconds), low power consumption, and exceptionally high dynamic range (HDR), making them particularly well suited for fast motion capture and extreme lighting conditions [31, 7, 47, 45, 34].
To better interpret event data, a fundamental task in event-based vision is event-to-video (E2V) reconstruction, which aims to recover intensity videos from raw event streams. A wide range of learning-based approaches have been proposed, typically adopting recurrent encoder–decoder architectures that aggregate events over time and decode latent representations into intensity frames [31, 33, 36, 8, 32]. Despite steady progress, most prior works largely frame E2V as a temporal–spatial signal recovery problem, implicitly assuming that object-level structure and scene context can be recovered from event streams through sufficiently strong temporal aggregation and feature learning (Fig. 1-b). However, due to the change-driven sensing mechanism, event signals are inherently semantically under-determined: they lack rich semantic cues (Fig. 1-a), such as object identity, texture, and global contextual information, that are naturally available in intensity frames [37, 9]. As a result, existing approaches that rely solely on temporal and spatial modeling learn event feature representations that are inherently poor in semantic content and attention (Fig. 1-d-left). Consequently, the reconstructed videos often remain semantically ambiguous at the object level, exhibiting distorted or inconsistent semantic details—e.g., inaccurate object boundaries or structural inconsistencies—especially in regions with weak or missing event evidence (Fig. 1-d-right). In contrast, semantic information plays a critical role in video reconstruction and scene understanding, as it provides high-level structural constraints and contextual priors that help disambiguate the ill-posed mapping from sparse observations to dense intensity signals [40, 21, 35].
From this perspective, we revisit the E2V task. Rather than viewing E2V solely as a process of temporally aggregating sparse events and extracting spatial features via a simple encoder-decoder architecture, we argue that effective reconstruction requires explicitly modeling the missing semantic information inherent in event signals. Accordingly, we propose a semantic-enriched end-to-end E2V training and supervision framework, called Semantic-E2VID, with three key focuses: semantic abstraction, semantic feature fusion, and semantic-aware supervision (Fig. 1-c). Our approach enables the model to learn high-level semantic cues that are not explicitly captured by event sensors, integrate these cues into semantically enriched event representations (Fig. 1-e-left), and leverage semantic-aware supervision to guide reconstruction toward structurally and semantically consistent video frames, even under sparse or ambiguous event conditions (Fig. 1-e-right). Specifically, to enable semantic abstraction for event data, we extract semantic knowledge from a pretrained vision foundation model, namely the Segment Anything Model (SAM) [13], through a Cross-modal Feature Alignment (CFA) module. This design avoids directly distilling frame-level representations into the event encoder, which can otherwise introduce modality-induced feature drift, and instead transfers semantic knowledge in a manner compatible with the event representation space. To effectively exploit the learned semantics during reconstruction, we further integrate semantic cues into event feature representations via a Semantic-aware Feature Fusion (SFF) block. This fusion strategy enables the decoder to jointly leverage temporal event information and semantic structure, allowing for a more faithful recovery of object boundaries and scene layout. In addition, we introduce a Semantic Perceptual E2V Supervision that explicitly emphasizes semantic consistency during training by guiding the reconstruction process toward object-centric regions. This supervision complements existing pixel-level and perceptual objectives and further reinforces the utilization of semantic cues during decoding.
Extensive experiments are conducted on a total of six public benchmarks, including three datasets with referenced ground-truth frames, two datasets with non-referenced frames, and one colorful reconstruction dataset. Semantic-E2VID consistently achieves superior performance over previous state-of-the-art E2V approaches across multiple benchmarks and evaluation metrics, demonstrating clear advantages in both quantitative accuracy and qualitative reconstruction quality. The key technical contributions of this work are summarized as follows:
  • 1.
    We revisit event-to-video reconstruction from a semantic perspective and propose Semantic-E2VID, a semantic-enriched E2V framework that explicitly addresses the semantic under-determinacy of event signals, going beyond conventional temporal and spatial modeling.
  • 2.
    We introduce principled semantic modeling mechanisms for E2V that enable semantic abstraction, semantic feature fusion, and semantic-aware supervision, allowing event representation to capture object-level structure and contextual information lost in the sensing process.
  • 3.
    Extensive experiments on six public benchmarks demonstrate that explicitly modeling and leveraging semantic information leads to more faithful reconstruction of object structure, clearer semantic boundaries, and improved robustness under challenging conditions.
2 Related Work
@@
Event-To-Video Reconstruction: Event cameras capture scene dynamics by recording asynchronous brightness changes, offering advantages such as low latency, high dynamic range, and energy efficiency. One fundamental task in event-based vision is event-to-video (E2V) reconstruction, where the goal is to synthesize intensity frames from raw event streams. works [31, 33, 1, 36, 8, 44, 11] have been done using either a recurrent-based, transformer-based, or Mamba-based network to learn event features over time for frame synthesis. More recent works [18, 53] leverage diffusion-based methods to tackle this problem; however, the time-consuming nature of diffusion models results in a heavy and inefficient forward process. Some [20, 54] also explored HDR frame reconstruction. Despite these advancements, most existing E2V methods overlook semantic information such as object categories, texture, relationships, positions, etc, which are missing in the event modality yet important for video reconstruction. [2] leverages text description for semantic-aware E2V, but introduces text, as an extra input modality. Our proposed Semantic-E2VID is the first work that addresses this gap by distilling visual semantic knowledge from SAM to improve semantic learning in event modality and reconstruction quality.
@@
Learning Semantics From SAM: The Segment Anything Model (SAM) [13] has emerged as a powerful vision foundation model, demonstrating strong generalization capabilities across diverse vision tasks [42, 12]. SAM’s success is attributed to its extensive pretraining on large-scale datasets, enabling it to produce robust and transferable semantic representations [41, 10]. Due to its strong semantic understanding, researchers have explored SAM for different downstream tasks. Works [49, 23, 5, 17] have first been done to transfer knowledge from SAM to medical image segmentation. Beyond that, some have utilized SAM as a teacher model for semantic knowledge distillation on other computer vision tasks, such as 2D Image Restoration [50] and feature learning and matching [46]. [3] distill knowledge from SAM for event-based segmentation. However, since both the teacher and student operate on the same task (segmentation), the distillation can be well aligned. In contrast, for the E2V task, directly distilling from frame-based models is challenging in transferring semantic knowledge from the frame modality to the event modality, as the objectives and data representations differ significantly [25].
3 Methodology
3.1 Overview
This section presents our semantic-aware event-to-video reconstruction framework. We begin by formulating the event-to-video task and introducing the input representation in Sec. 3.2. We then discuss the established non-semantic event-to-video reconstruction baseline that serves as the spatiotemporal backbone in Sec. 3.3. Building upon this backbone, we introduce Semantic-E2VID in Sec. 3.4, which explicitly injects high-level semantic abstraction into event representations through semantic learning and feature fusion. Finally, we present the semantic-aware supervision strategy in our end-to-end framework in Sec. 3.5.
3.2 Modality and Task Formulation
@@
where $E$ represents the recorded events, while $I_{t}$ and $I_{t-w}$ denote the intensity images at two different timestamps with a time window $w$. The function $\Gamma\left\{\theta,\epsilon\right\}$ maps log-intensity changes to discrete event outputs, with $c$ serving as an offset to avoid $\log(0)$. The parameter $\epsilon$ defines the event triggering threshold. Specifically, $\Gamma\{\theta,\epsilon\}=1$ when $\theta\geq\epsilon$, meaning a positive event triggered, and $\Gamma\{\theta,\epsilon\}=-1$ when $\theta\leq-\epsilon$, meaning a negative event triggered [6, 48, 22].
@@
3.3 Spatiotemporal E2V Backbone
We first establish a non-semantic event-to-video reconstruction baseline that serves as a temporal-spatial backbone by following prior works [31, 36]. Specifically, we adopt a U-Net-based recurrent encoder–decoder architecture originally proposed in [31], which serves as a representative backbone for modeling temporal dynamics and spatial structures in event streams. The encoder consists of a sequence of convolutional layers and ConvLSTM units, where each convolutional layer uses a kernel size of $5$ with a stride $2$, and each ConvLSTM employs a kernel size of $3$ to preserve spatial and channel consistency between inputs and hidden states. Residual blocks are constructed using two convolutional layers with a kernel size $3$. Given an input event voxel grid $V_{k}\in\mathbb{R}^{W\times H\times B}$, the encoder extracts event features $\mathcal{F}_{e}\in\mathbb{R}^{\frac{1}{2}W\times\frac{1}{2}H\times 256}$. The decoder mirrors the encoder structure and is composed of bilinear upsampling followed by convolutional layers, progressively reconstructing intensity frames from the encoded event features. Additionally, following [8], the decoder incorporates auxiliary context fusion mechanisms, which dynamically generate convolutional weights conditioned on both current event inputs and previously reconstructed frames to enhance temporal coherence.
Despite their effectiveness, such architectures fundamentally focus on temporal and spatial modeling of event data and do not explicitly address the semantic under-determinacy inherent in event signals. As a result, the learned event representations remain limited in semantic content, making it difficult to recover object-level structure and contextual information from sparse event observations.
@@ 3.3 Spatiotemporal E2V Backbone
Figure 2: Overview of the proposed Semantic-E2VID framework. Given input events, an encoder first extracts event features $\mathcal{F}_{e}$. To enable semantic abstraction without introducing modality-induced feature drift, we introduce the Cross-Modal Feature Alignment (CFA) module that aligns $\mathcal{F}_{e}$ with frame semantic features $\mathcal{F}_{sam}$ extracted from a pretrained vision foundation model (SAM), yielding learned frame semantic representations $\mathcal{F}_{semantic}$. To effectively inject semantic knowledge into the event representation space, the Semantic-Aware Feature Fusion (SFF) block integrates $\mathcal{F}_{semantic}$ with $\mathcal{F}_{e}$, producing semantically enriched event features $\mathcal{F}_{e}^{\prime}$. A decoder then reconstructs the final intensity frame from $\mathcal{F}_{e}^{\prime}$. During training, the reconstruction process is jointly supervised by conventional reconstruction and temporal losses, together with semantic-aware supervisions that emphasize semantic consistency in object-centric regions.
3.4 Semantic-E2VID
Motivated by this limitation, we revisit the E2V task from a semantic perspective. Rather than treating reconstruction purely as a temporal-spatial signal recovery problem, we reformulate E2V as a semantic-enriched reconstruction task, where effective reconstruction requires explicitly learning, fusing, and supervising high-level semantic information that is absent in event data. In particular, effective semantic-enriched E2V necessitates three key capabilities: (i) semantic abstraction, i.e., extracting high-level semantic knowledge from complementary modalities and aligning it with event representations; (ii) semantic feature fusion, i.e., integrating learned semantics with encoded event features in a manner compatible with the event representation space; and (iii) semantic-aware supervision, i.e., guiding the reconstruction process toward object-level semantic consistency.
To realize these capabilities in an end-to-end manner, we propose Semantic-E2VID, a semantic-guided E2V framework composed of three corresponding components: Cross-Modal Feature Alignment (CFA) module for semantic abstraction, Semantic-aware Feature Fusion (SFF) block for injecting semantics into event features, and Semantic Perceptual E2V Supervision strategy for enforcing semantic consistency during reconstruction. An overview of the framework is shown in Fig. 2, and each component is detailed in the following subsections.
3.4.1 Semantic Abstraction and Learning.